A discrete random variable has pmf P(X=n)=(1/2)^n. Let Y={1, if x is even and -1 if x is odd

Find the expected value of Y.
1
Expert's answer
2020-04-27T19:20:27-0400

"P(X=n)=(1\/2)^n"

"Y=\begin{cases}
1, \ \text{if x is even}
\\
-1, \ \text{if x is odd}
\end{cases}"

"P(Y=1)=\sum \limits_{n=1}^{\infin}P(X=2n)= \sum \limits_{n=1}^{\infin} (1/2)^{2n}= \sum \limits_{n=1}^{\infin} (1/4)^{n}=\frac{1/4}{1-1/4}=1/3"

"P(Y=-1)=1-P(Y=1)=1-1\/3=2\/3"

"E[Y]=1\times P(Y=1)+(-1)\times P(Y=-1)=1/3-2/3=-1/3"

Answer: the expected value of "Y" is -1/3.